OpenStudy (anonymous):

If f(x) = 5x + 3, then what is f–1(x) ?

OpenStudy (anonymous):

inverse

OpenStudy (anonymous):

f(x)=y=5x+3 sfr x and then swap x and y

OpenStudy (anonymous):

interchange x and y and solve for y

OpenStudy (anonymous):

y=5x+3 y-5x=3 -5x=3-y x=-3/5 - y/3 so y=-3/5 - x/3

OpenStudy (anonymous):

I got x+3/5 as my answer?

OpenStudy (anonymous):

\[y=-3/5+1/5\,x\]

OpenStudy (anonymous):

I'll draw the options

OpenStudy (anonymous):